OpenStudy (anonymous):

Use the shell method to find the volume between y=e^x , y=e^-x nad x=1 about the y -axis

OpenStudy (anonymous):

i got 4,623 as my answer. i can walk you through what i did if my answer is wrong

OpenStudy (anonymous):

anyone?

OpenStudy (phi):

the set-up should be \[ 2 \pi \int_0^1 (e^x - e^{-x}) x \ dx \]

OpenStudy (anonymous):

i've got that, i then split it into 2 integrals; Sxe^x-Sxe^-x and used integration by parts to solve

OpenStudy (phi):

yes. sounds correct.

OpenStudy (anonymous):

this left me with \[2\pi[(x e^{x}-e^{x})-(xe^{-x}-e^{-x})]\] from 0 to 1

OpenStudy (phi):

I get \[ 2\pi[(x e^{x}-e^{x})-(-xe^{-x}-e^{-x}) ]\bigg|_0^1 \]

OpenStudy (anonymous):

looks like i forgot to type the -, i got the same

OpenStudy (phi):

the final answer is 4 pi / e

OpenStudy (anonymous):

i didnt write it in that form, but i got the same 4.623

OpenStudy (phi):

yes. But it is probably best to write it in exact form. Then the final decimal version can be rounded to the appropriate precision, depending on how it is being used. Also, 4 pi /e looks nicer

OpenStudy (anonymous):

will do, thanks!